1/3(p+1)is greater than or equal to 3/6(p-2)

Answer by Alan3354(69443)

1/3(p+1)is greater than or equal to 3/6(p-2)
-----
(p+1)/3 >= (p-2)/2
Multiply by 6
2(p+1) >= 3(p-2)
2p+2 >= 3p-6
8 >= p
p <= 8